OpenStudy (anonymous):

what percent of 240 is 30

OpenStudy (anonymous):

240/30 = 0.125 (12.5%)

OpenStudy (anonymous):

i Mean 30/240 = 0.125 (12.5%)

OpenStudy (anonymous):

what is the difference

OpenStudy (anonymous):

I wrote the first one wrong. it's 30/240 not 240/30

OpenStudy (anonymous):

When something is a percentage of something else, you're basically saying: \[\LARGE \frac{x}{100}=\frac{30}{240}\] Solve for X.